Solve by elimination
4x+5y=-16
5x-6y=-20

Answer by praseena(37)

4x+5y=-16
5x-6y=-20
-----------------------------------------------------------------------
Multiply the first equation by 6 and the second equation by 5
6(4x+5y)=6*-16
6*4x+6*5y=-96
24x+30y=-96 -------(a)
Multiply the second equation by 5,
5 (5x-6y) =5*20
5*5x - 5 * 6y = -100
25x - 30y = -100 ------(b)
---------------------------------------------------
Add equations (a) and (b)
===> 24x - 30y + 25x + 30y = -96 + - 100
===> 24x + 25y + -30y + 30y = -196
===> 49x = -196
Divide on both sides by 49
===> x = -196/49 = -4
----------------------------------------------------------------
Now substitute the value x=-4 in any one of the given pair of equations.
==> 4 * -4 + 5y = -16
==> -16 + 5y = -16
Add 16 on both sides
==> 5y = 0
Divide both sides by 5
==> y=0
=========================================================================
So the solution for the given pair of equations is,x=-4 and x=0.
Bye